OpenStudy (anonymous):

( 6 )^3 * ( x^6 )^5 * ( w^6 )^1 * ( x^9 )^5

OpenStudy (anonymous):

Please help me understand step by step not just the answer thanks 😄

OpenStudy (jdoe0001):

\(\bf ( 6 )^3 \times ( x^6 )^5 \times ( w^6 )^1 \times ( x^9 )^5\implies ( 6^1 )^3 \times ( x^6 )^5 \times ( w^6 )^1 \times ( x^9 )^5\\ \quad \\ \Large 6^{1\times 3}\quad \times \quad x^{6\times 5}\quad \times \quad w^{6\times 1}\quad \times \quad x^{9\times 5}\)

OpenStudy (shamil98):

when you multiply exponents you add them @jdoe0001

OpenStudy (anonymous):

Thanks @jdoe0001

OpenStudy (jdoe0001):

well, yes IF IT'S THE SAME BASE, that's true for the "x" one like \(\bf 6^3\times x^{30}\times w^6\times x^{45}\implies 6^3\times (x^{30}\times x^{45})\times w^6\implies 6^3\times (x^{30+45})\times w^6\\ \quad \\ 6^3\times x^{75}\times w^6\implies 6^3 x^{75}w^6\)

OpenStudy (jdoe0001):

yw